# Cosmology in Modified $f(\mathcal{G})$ Gravity: A Late Time Cosmic Phenomena

### v1

@@ Abstract
Abstract: In this paper, we have introduced a numerical method for solving the Friedmann equations of modified $f(\mathcal{G})$ gravity in the presence of pressureless matter, enabling us to predict the redshift behavior of the Hubble expansion rate. To evaluate the credibility of the model, we applied a Bayesian MCMC technique using late-time cosmic observations to impose limitations on the free parameters of the theory. Our results suggest that the $f(\mathcal{G})$ model can reproduce the low-redshift behavior of the standard $\Lambda$CDM model, but there are significant differences at high redshifts, leading to the absence of a standard matter-dominated epoch. We also examined the profiles of cosmographic parameters using the model parameter values from the standard range to verify the intermediate epochs. Our analysis shows that the highly promising $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs. We presented a dynamical system analysis framework to examine the stability of the model. Our study identified critical points depicting various phases of the Universe and explained the evolutionary epochs. We demonstrated that the model effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternate explanation for the observed acceleration of the Universe.
Keywords: Gauss-Bonnet Invariant, Cosmological Datasets, Dynamical System Analysis.
I Introduction
@@
The quest to understand the gravitational force led to the development of General Relativity (GR) by Albert Einstein in 1915 [1]. This theory revolutionized our comprehension of gravity by describing it as a curvature of spacetime caused by the presence of mass and energy [2]. While GR has successfully passed numerous experimental tests [3] and remains the cornerstone of modern gravitational physics, it struggles to fully explain phenomena such as the accelerated expansion of the universe [4], dark matter [5], and dark energy (DE) [6]. In response to these challenges, several modified theories of gravity have been proposed [7].
@@
The most recent observational data, including the Dark Energy Spectroscopic Instrument (DESI) surveys [8], Type Ia Supernovae (SNe Ia) [4, 9], Wilkinson microwave anisotropy probe experiment (WMAP) [10], cosmic microwave background (CMB) [11], Baryon oscillation spectroscopic survey (BOSS) [12] and the Baryon Acoustic Oscillations (BAO) datasets [13] have prompted researchers to consider modifications and expansions to the principles of GR in theories such as $f(R)$ [14], $f(T)$ [15], and $f(Q)$ [16, 17]. These alternative theories aim to better accommodate and explain the new observational data. One of the most straightforward extensions to Einstein’s gravity is the so-called $f(R)$ gravity, where $f$ is an arbitrary function of the Ricci scalar $R$ [18]. Even in this relatively simple case, constructing viable $f(R)$ models consistent with cosmological and local gravity constraints is not straightforward. This complexity arises because $f(R)$ gravity introduces a strong coupling between DE and non-relativistic matter in the Einstein frame [19]. Extensions of GR, including the Gauss-Bonnet (GB) invariant in the gravitational action, have generated significant interest [20, 21, 22, 23, 24, 25]. One such theory that has garnered significant interest is $f(\mathcal{G})$ gravity. This theory modifies the Einstein-Hilbert action by introducing a function of the GB invariant, denoted as $\mathcal{G}$, which is a combination of the Ricci scalar $R$, the Ricci tensor $R_{\mu\nu}$, and the Riemann tensor $R_{\mu\nu\sigma\gamma}$ [26, 27, 28, 20, 29]. It belongs to an infinite class of curvature invariants known as the Lovelock scalars along with $R$. These do not introduce derivative terms higher than two into the equations of motion for the metric tensor. In four dimensions, the term $\sqrt{-g}\mathcal{G}$ is a total derivative, so the GB term contributes to the equations of motion only when coupled to something else, such as a scalar field $\phi$ with the form $f(\phi)\mathcal{G}$ coupling [30, 31]. A dilaton-graviton mixing term generates this kind of coupling in the low-energy effective action of string theory [32].
@@
The interest in $f(\mathcal{G})$ gravity lies in its potential to explain the observed late-time cosmic acceleration in the Universe. This acceleration could be caused by a gravity modification rather than an unusual matter source with negative pressure [33]. In recent years, significant research has been conducted into modified gravity to understand the nature of DE [34]. Modified gravity models are particularly attractive because they align more closely with cosmological observations and local gravity experiments than models that rely on exotic matter sources [35]. It is suggested that this theory can pass solar system tests [36, 23] and may describe the most exciting features of late-time cosmology, such as the transition from deceleration to acceleration and the current acceleration of the universe [20, 36, 23, 37].
@@
In this work, we explored a subclass of the $f(\mathcal{G})$ model to test its viability as an alternative to the standard cosmological paradigm. We have developed a numerical method to predict the redshift behavior of the Hubble expansion rate, and our results suggest that the model can reproduce the low-redshift behavior of the standard $\Lambda$CDM model but has significant differences at high redshifts. The $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs and effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternate explanation for the observed acceleration of the Universe. We delved into the background cosmological dynamics of the chosen model and evaluated its feasibility using Bayesian analysis supported by MCMC methods applied to late-time cosmic observations, such as Supernovae Ia (Pantheon+) and observational Hubble data (CC sample). We introduced a dynamical system analysis framework to assess the stability of the model. Our research pinpointed critical points that illustrate different phases of the Universe and elucidated the evolutionary epochs. We showed that the model effectively represents the changing energy components over cosmic time, supporting its credibility as an alternative explanation for the observed acceleration of the Universe.
@@
This work aims to establish constraints on $f(\mathcal{G})$ cosmology models using CC and Pantheon+ datasets. The paper comprehensively analyses $f(\mathcal{G})$ gravity and uses dynamical system analysis to investigate the stability of the model. The mathematical formalism of $f(\mathcal{G})$ gravity is detailed in Section II. In Section III, we utilize the Markov Chain Monte Carlo (MCMC) method to establish correlations between the $f(\mathcal{G})$ gravity model and observational data to determine the best fits for the model parameters $H_{0}$, $\alpha$, $\beta$, and $m$. Additionally, we present plots of various cosmological parameters such as deceleration, effective equation of state (EoS) $\omega=\frac{p}{\rho}$, Om diagnostic, and the $r-s$ parameter plot, which are essential for understanding the dynamical behavior of the Universe under $f(\mathcal{G})$ gravity. Subsequently, in Section IV, we construct a dynamical system framework to analyze the critical points of the $f(\mathcal{G})$ gravity model. This analysis is crucial for assessing the stability and viability of the model and its alignment with current cosmological observations. Finally, in Section V, we present the conclusions of our results.
II $R+f(\mathcal{G})$ gravity
We consider the following action, which describes GR plus a function of the GB term [20, 38]:
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa^{2}}R+f(\mathcal{G})+L_{m}\right],$ (1)
@@ II $R+f(\mathcal{G})$ gravity
where $\kappa^{2}=8\pi G_{N}$, $G_{N}$ being the Newtonian constant, and the GB invariant is defined as usual:
@@
where we made the notations $f_{\mathcal{G}}=\frac{df}{d\mathcal{G}}$ and $f_{\mathcal{G}\mathcal{G}}=\frac{d^{2}f}{d\mathcal{G}^{2}}$. We shall assume throughout the paper a spatially flat FLRW Universe, whose metric is given by
@@
where $a(t)$ is the scale factor at cosmological time $t$. For the metric (4), the field equations give the FLRW equations, with the form
00 0 $=$ $-\frac{3}{\kappa^{2}}H^{2}+\mathcal{G}f_{\mathcal{G}}-f(\mathcal{G})-24\dot{\mathcal{G}}H^{3}f_{\mathcal{G}\mathcal{G}}+\rho_{m},$ (5a)
00 0 $=$ $8H^{2}\ddot{f}_{\mathcal{G}}+16H\left(\dot{H}+H^{2}\right)\dot{f}_{\mathcal{G}}+\frac{1}{\kappa^{2}}\left(2\dot{H}+3H^{2}\right)$ (5b)
$+f-\mathcal{G}f_{\mathcal{G}}+p_{m}.$
The Hubble rate $H$ is here defined by $H=\dot{a}/a$, while the matter energy density $\rho_{m}$ satisfies the standard continuity equation:
$\dot{\rho}_{m}+3H(1+w)\rho_{m}=0,$ (6)
and the GB invariant $\mathcal{G}$ and the Ricci scalar $R$ can be defined as functions of the Hubble parameter as
@@
II.1 Power-Law $f(\mathcal{G})$ Model
In light of the multiple cosmological data analyses and tests conducted within our solar system, all confirming the principles of GR, we can conclude that any departures from standard GR are expected to be negligible. Consequently, we are led to consider the following $f(\mathcal{G})$ functional [36, 23],
$f(\mathcal{G})=\alpha\sqrt{\beta}\left(\frac{\mathcal{G}^{2}}{\beta^{2}}\right)^{m}$ (8)
@@ II.1 Power-Law $f(\mathcal{G})$ Model
In order to determine the theoretical values of the Hubble rate, we can calculate it by solving equation (5a) numerically. When we assume that matter behaves as a pressureless perfect fluid, we can express the matter density as $\rho_{m}=3H_{0}^{2}\Omega_{m0}(1+z)^{2}$, with $\Omega_{m0}$ representing the current value of the matter density parameter. Thus, for the particular model we are examining, the first Friedmann equation can be expressed as follows:
@@
Equation (II.1) represents a second-order differential equation for the function $H(z)$, which can be solved using appropriate boundary conditions. The first initial condition is simply $H(0)=H_{0}$. For the second initial condition to be determined, we can ensure that, at present, the first derivative of the Hubble parameter is consistent with the predictions of the standard $\Lambda$CDM model, which is characterized by the following expansion law:
$H_{\Lambda\text{CDM}}=H_{0}\sqrt{1-\Omega_{m0}+\Omega_{m0}(1+z)^{3}}$ (10)
After differentiating the above equation with respect to $z$, the second initial condition for Eq. (II.1) is obtained as $H^{\prime}(0)=\frac{3H_{0}\Omega_{m0}}{2}$.
III Observation with numerical solution
In this part, we will assess the observational feasibility of the model under examination by conducting a Bayesian analysis of the late-time cosmic data. Specifically, we will evaluate the data from the SNe Ia Pantheon+ sample [39] and the cosmic chronometers (CC) derived from the observational Hubble data compiled in Ref. [40]. Utilizing these datasets for statistical analyses enables us to obtain dependable results that are not influenced by assumptions of any particular underlying reference model [41, 42, 43]. In the following subsections, we will outline the key characteristics of these measurements and the corresponding Likelihood functions.
III.1 Cosmic Chronometers (CC)
@@
$H(z)=\frac{\dot{a}}{a}=-\frac{1}{1+z}\frac{dz}{dt}\approx-\frac{1}{1+z}\frac{\Delta z}{\Delta t}\,$ (11)
Here, $\dot{a}$ is the derivative of the scale factor $a$ with respect to time $t$, and $\Delta z$ and $\Delta t$ are the differences in redshift and time, respectively, between two objects. The value of $\Delta z$ can be determined via a spectroscopic survey, while the differential ages $\Delta t$ of passively evolving galaxies can be used to estimate the value of $H(z)$. Compiling such observations can be regarded as a CC sample. We use 32 objects covering the redshift range $0.07\leq z\leq 1.965$ [40, 43]. For these measurements, one can construct a $\chi^{2}_{\text{CC}}$ estimator as follows:
$\chi^{2}_{CC}=\sum_{i=1}^{32}\frac{[H_{\text{th}}(z_{i})-H_{\text{obs}}(z_{i})]^{2}}{\sigma^{2}_{H}(z_{i})}\,.$ (12)
@@
III.2 Supernovae type Ia (SNe Ia)
We will also take into account the Pantheon+ SNe Ia dataset, which includes 1701 measurements of the relative luminosity distance of SNe Ia spanning the redshift range of $2.26130.00122<z<$ 2.2613 [39]. The Pantheon+ compilation consists of distance moduli derived from 1701 light curves of 1550 spectroscopically confirmed SNe Ia within the redshift range of $2.26130.00122<z<$ 2.2613, collected from 18 different surveys. It is worth noting that 77 of the 1701 light curves are associated with galaxies containing Cepheids. The Pantheon+ dataset is advantageous in that it can also be used to limit the value of $H_{0}$ besides the model parameters. To estimate the model parameter from the Pantheon+ samples, we minimize the $\chi^{2}$ function. To calculate the chi-square $(\chi^{2}_{\text{SNe}})$ value using the Pantheon compilation of 1701 supernovae data points, we use the following formula:
$\chi^{2}_{SNe}=\Delta\mu^{T}(C_{Sys+Stat}^{-1})\Delta\mu\,,$ (13)
where $C_{\text{Sys}+\text{Stat}}^{-1}$ is the covariance matrix of the Pantheon+ dataset, which includes systematic and statistical uncertainties. As defined below, $\Delta\mu$ denotes the distance residual:
$\Delta\mu_{k}=\mu_{k}-\mu_{th}(z_{k})\,.$ (14)
The equation above defines $\mu_{k}$ as the distance modulus of the $k^{th}$ Type Ia supernova. It is important to note that $\mu_{k}$ is calculated as the difference between the apparent magnitude of the $k^{th}$ SNe, denoted by $m_{Bk}$, and the fiducial magnitude of an SNe, denoted by $M$.
The theoretical distance modulus $\mu_{th}$ is calculated using the expression
$\mu_{th}(z,\vartheta)=5\log_{10}\left(\frac{d_{L}(z,\vartheta)}{1\text{ Mpc}}\right)+25\,,$ (15)
where $d_{L}(z,\vartheta)$ represents the luminosity distance in Mpc depending on the model, which is
$d_{L}(z,\vartheta)=\frac{c(1+z)}{H_{0}}\int_{0}^{z}\frac{d\zeta}{E(\zeta)}\,,$ (16)
where $E(z)=\frac{H(z)}{H_{0}}$ and $c$ stands for the speed of light. Additionally, the residual distance is denoted by
$\Delta\bar{\mu}=\begin{cases}\mu_{k}-\mu_{k}^{cd},&\text{if $k$ is in Cepheid hosts}\\ \mu_{k}-\mu_{th}(z_{k}),&\text{otherwise}\end{cases}$ (17)
where $\mu_{k}^{cd}$ is the Cepheid host-galaxy distance that SH0ES revealed. This covariance matrix can be coupled with the SNe covariance matrix to construct the covariance matrix for the Cepheid host galaxy. Equipped with statistical and systematic uncertainties from the Pantheon+ dataset, the combined covariance matrix is represented as $C^{SNe}_{Sys+Stat}+C^{cd}_{Sys+Stat}$. This expression defines the $\chi^{2}$ function for the combined covariance matrix used to constrain cosmological models in the analysis:

### v2

# MNRAS  template – title goes here

@@ Abstract
In this work, we present a method for numerically solving the Friedmann equations of modified $f(\mathcal{G})$ gravity in the presence of pressureless matter. This method enables us to predict the redshift behavior of the Hubble expansion rate. To evaluate the credibility of the model, we applied a Bayesian MCMC technique using late-time cosmic observations to impose limitations on the free parameters of the Gauss-Bonnet model. Our results suggest that the $f(\mathcal{G})$ model can reproduce the low-redshift behavior of the standard $\Lambda$CDM model, but there are significant differences at high redshifts, leading to the absence of a standard matter-dominated epoch. We also examined the profiles of cosmographic parameters using the model parameter values from the standard range to verify the intermediate epochs. Our analysis shows that the highly promising $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs. We presented a dynamical system analysis framework to examine the stability of the model. Our study identified critical points depicting various phases of the Universe and explained the evolutionary epochs. We demonstrated that the model effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternate explanation for the observed acceleration of the Universe.
keywords:
cosmology: theory, cosmology: dark energy
pubyear: 2024pagerange: MNRAS  template – title goes hereReferences
1 Introduction
@@
The quest to understand the gravitational force led to the development of General Relativity (GR) by Einstein (1915). This theory revolutionized our understanding of gravity by describing it as the curvature of spacetime caused by the presence of mass and energy (Misner et al., 1973). Although GR has successfully passed numerous experimental tests (Will, 2014) and remains the cornerstone of modern gravitational physics, it struggles to fully explain phenomena such as the accelerated expansion of the Universe (Riess et al., 1998), dark matter (Bertone et al., 2005), and dark energy (DE) (Copeland et al., 2006). In response to these challenges, several modified theories of gravity have been proposed (Clifton et al., 2012).
@@
The most recent observational data, including the Dark Energy Spectroscopic Instrument (DESI) surveys (Adame et al., 2024), Type Ia Supernovae (SNe Ia) (Riess et al., 1998; Perlmutter et al., 1999), Wilkinson microwave anisotropy probe experiment (WMAP) (Spergel et al., 2003), cosmic microwave background (CMB) (Hinshaw et al., 2013), Baryon oscillation spectroscopic survey (BOSS) (Alam et al., 2017) and the Baryon Acoustic Oscillations (BAO) datasets (Eisenstein et al., 2005) have prompted researchers to consider modifications and expansions to the principles of GR in theories such as $f(R)$ (Sotiriou & Faraoni, 2010), $f(T)$ (Ferraro & Fiorini, 2007), and $f(Q)$ (Jiménez et al., 2018; Heisenberg, 2023). These alternative theories aim to better accommodate and explain the new observational data. One of the most straightforward extensions to Einstein’s gravity is the so-called $f(R)$ gravity, where $f$ is an arbitrary function of the Ricci scalar $R$ (Nojiri & Odintsov, 2007). Even in this relatively simple case, constructing viable $f(R)$ models consistent with cosmological and local gravity constraints is not straightforward. This complexity arises because $f(R)$ gravity introduces a strong coupling between DE and non-relativistic matter in the Einstein frame (Capozziello & Francaviglia, 2008). Extensions of GR, including the Gauss-Bonnet (GB) invariant in the gravitational action, have generated significant interest (Nojiri & Odintsov, 2005; Li et al., 2007; Elizalde et al., 2010; De Felice & Tsujikawa, 2009; Maurya et al., 2021, 2022; Nojiri et al., 2017; Odintsov et al., 2019; Lohakare et al., 2023; Nojiri et al., 2008). One such theory that has garnered significant interest is $f(\mathcal{G})$ gravity. This theory modifies the Einstein-Hilbert action by introducing a function of the GB invariant, denoted $\mathcal{G}$, which is a combination of the Ricci scalar $R$, the Ricci tensor $R_{\mu\nu}$, and the Riemann tensor $R_{\mu\nu\sigma\gamma}$ (Stelle, 1978; Barth & Christensen, 1983; Felice & Suyama, 2009; Nojiri & Odintsov, 2005, 2006). It belongs to an infinite class of curvature invariants known as the Lovelock scalars along with $R$. These do not introduce derivative terms greater than two into the equations of motion for the metric tensor. In four dimensions, the term $\sqrt{-g}\mathcal{G}$ is a total derivative, so the GB term contributes to the equations of motion only when coupled to something else, such as a scalar field $\phi$ with the form $f(\phi)\mathcal{G}$ coupling (Tsujikawa et al., 2002; Cartier et al., 2000). A dilaton-graviton mixing term generates this kind of coupling in the low-energy effective action of string theory (Gasperini & Veneziano, 2003).
@@
The interest in $f(\mathcal{G})$ gravity lies in its potential to explain the observed late-time cosmic acceleration in the Universe. This acceleration could be caused by a gravity modification rather than an unusual source of matter with negative pressure (Carroll & other, 2004). In recent years, significant research has been conducted into modified gravity to understand the nature of DE (Nojiri & Odintsov, 2011). Modified gravity models are particularly attractive because they align more closely with cosmological observations and local gravity experiments than models that rely on exotic matter sources (Joyce et al., 2015). It is suggested that this theory can pass solar system tests (Davis, 2007; De Felice & Tsujikawa, 2009) and may describe the most exciting features of late-time cosmology, such as the transition from deceleration to acceleration and the current acceleration of the Universe (Nojiri & Odintsov, 2005; Davis, 2007; De Felice & Tsujikawa, 2009; Cognola et al., 2006).
@@
In this work, we have explored a subclass of the $f(\mathcal{G})$ model to test its viability as an alternative to the standard cosmological paradigm. We have developed a numerical method to predict the redshift behavior of the Hubble expansion rate, and our results suggest that the model can reproduce the low-redshift behavior of the standard $\Lambda$CDM model but has significant differences at high redshifts. The $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs and effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternative explanation for the observed acceleration of the Universe. We delved into the background cosmological dynamics of the chosen model and evaluated its feasibility using Bayesian analysis supported by Markov Chain Monte Carlo (MCMC) methods applied to late-time cosmic observations, such as Supernovae Ia (Pantheon+) and observational Hubble data (CC sample). We have introduced a dynamical system analysis framework to assess the stability of the model. Our research pinpointed critical points that illustrate different phases of the Universe and elucidated the evolutionary epochs. We have shown that the model effectively represents the changing energy components over cosmic time, supporting its credibility as an alternative explanation for the observed acceleration of the Universe.
@@
This work aims to establish constraints on $f(\mathcal{G})$ cosmology models using CC and Pantheon+ datasets. The paper comprehensively analyses $f(\mathcal{G})$ gravity and uses dynamical system analysis to investigate the stability of the model. The mathematical formalism of $f(\mathcal{G})$ gravity is detailed in Section 2. In Section 3, we use the MCMC method to establish correlations between the $f(\mathcal{G})$ gravity model and observational data to determine the best fits for the model parameters $H_{0}$, $\alpha$, $\beta$, and $m$. Additionally, we present plots of various cosmological parameters such as deceleration, effective equation of state (EoS) $\omega_{\text{eff}}=\frac{p_{\text{eff}}}{\rho_{\text{eff}}}$, Om diagnostic, and the $r-s$ parameter plot, which are essential for understanding the dynamical behavior of the Universe under $f(\mathcal{G})$ gravity. Subsequently, in Section 4, we construct a dynamical system framework to analyze the critical points of the $f(\mathcal{G})$ gravity model. This analysis is crucial for assessing the stability and viability of the model and its alignment with current cosmological observations. Finally, in Section 5, we present the conclusions of our results.
2 $R+f(\mathcal{G})$ gravity
We consider an action that encompasses GR and a functional dependent on the GB term (Nojiri & Odintsov, 2005; Nojiri et al., 2006):
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa^{2}}R+f(\mathcal{G})+\mathcal{L}_{m}\right],$ (1)
@@ 2 $R+f(\mathcal{G})$ gravity
where $\kappa^{2}=8\pi G_{N}=1$, $G_{N}$ being the Newtonian constant, $\mathcal{L}_{m}$ denotes the matter Lagrangian. The GB topological invariant is defined as:
@@
where we made the notations $f_{\mathcal{G}}=\frac{df}{d\mathcal{G}}$ and $f_{\mathcal{G}\mathcal{G}}=\frac{d^{2}f}{d\mathcal{G}^{2}}$. We assume a spatially flat FLRW Universe throughout the paper, with a metric given by
@@
where $a(t)$ represents the scale factor at cosmological time $t$ and the GB invariant $\mathcal{G}$ and the Ricci scalar $R$ can be defined as functions of the Hubble parameter as
@@
The field equations for the metric (4) yield the FLRW equations in the form:
$3H^{2}$ $=$ $\kappa^{2}\left(\rho_{\text{m}}+\rho_{\text{r}}+\rho_{\text{DE}}\right)=\kappa^{2}\rho_{\text{eff}},$ (6)
$\left(2\dot{H}+3H^{2}\right)$ $=$ $-\kappa^{2}\left(\frac{\rho_{\text{r}}}{3}+p_{\text{DE}}\right)=-\kappa^{2}p_{\text{eff}},$ (7)
where $\rho_{\text{m}}$, $\rho_{\text{r}}$ and $\rho_{\text{DE}}$ denotes the matter density, radiation density and DE density, respectively. The over-dot indicates a derivative with respect to cosmic time $t$, and $H\equiv\dot{a}/a$ represents the Hubble parameter. Furthermore, the effective DE density and pressure have been defined as follows:
$\rho_{\text{DE}}=\frac{1}{\kappa^{2}}\Big{[}\mathcal{G}f_{\mathcal{G}}-f(\mathcal{G})-24\dot{\mathcal{G}}H^{3}f_{\mathcal{G}\mathcal{G}}\Big{]},$ (8a)
$p_{\text{DE}}=\frac{1}{\kappa^{2}}\Big{[}8H^{2}\ddot{f}_{\mathcal{G}}+16H\left(\dot{H}+H^{2}\right)\dot{f}_{\mathcal{G}}+f-\mathcal{G}f_{\mathcal{G}}\Big{]},$ (8b)
Without interactions between non-relativistic matter and radiation, these components independently follow their respective conservation laws $\dot{\rho}_{\text{m}}+3H\rho_{\text{m}}=0$ and $\dot{\rho}_{\text{r}}+4H\rho_{\text{r}}=0$. From Equations (8a) and (8b), it can be concluded that the DE density and pressure follow the standard evolution equation:
$\dot{\rho}_{\text{DE}}+3H\left(\rho_{\text{DE}}+p_{\text{DE}}\right)=0.$ (9)
The parameter for the effective dark energy EoS can be defined as follows:
$\omega_{\text{DE}}=\frac{p_{\text{DE}}}{\rho_{\text{DE}}}.$ (10)
In the $\Lambda$CDM limit, as expected, the EoS parameter $\omega_{\text{DE}}\to-1$.
2.1 Power-Law $f(\mathcal{G})$ Model
In light of the multiple cosmological data analyses and tests conducted within our solar system, all confirming the principles of GR, one can conclude that any departures from standard GR are expected to be negligible. Consequently, we led to consider the following $f(\mathcal{G})$ functional (Davis, 2007; De Felice & Tsujikawa, 2009),
$f(\mathcal{G})=\alpha\sqrt{\beta}\left(\frac{\mathcal{G}^{2}}{\beta^{2}}\right)^{m}\,,$ (11)
@@ 2.1 Power-Law $f(\mathcal{G})$ Model
In order to determine the theoretical values of the Hubble rate, we can calculate it by solving equation (8a) numerically. When we assume that matter behaves as a pressureless perfect fluid, we can express the matter density as $\rho_{m}=3H_{0}^{2}\Omega_{\text{m}0}(1+z)^{3}$, with $z$ representing the cosmological redshift (defined as $\frac{a_{0}}{a}=1+z$, where $a_{0}$ represents the scale factor at present and $a$ denotes the scale factor when the light emitted) and $\Omega_{\text{m}0}$ representing the current value of the matter density parameter. Thus, for the particular model we are examining, the first Friedmann equation can be expressed as follows:
@@
Equation (2.1) represents a second-order differential equation for the function $H(z)$, which can be solved using appropriate boundary conditions. The first initial condition is simply $H(0)=H_{0}$. For the second initial condition to be determined, we can ensure that, at present, the first derivative of the Hubble parameter is consistent with the predictions of the standard $\Lambda$CDM model, which is characterized by the following expansion law:
$H_{\Lambda\text{CDM}}=H_{0}\sqrt{1-\Omega_{\text{m}0}+\Omega_{\text{m}0}(1+z)^{3}}\,,$ (13)
After differentiating the above equation with respect to $z$, the second initial condition for equation (2.1) is obtained as $H^{\prime}(0)=\frac{3}{2}H_{0}\Omega_{\text{m}0}$.
3 Observation with numerical solution
In this part, we will assess the observational feasibility of the model under examination by conducting a Bayesian analysis of the late-time cosmic data. Specifically, we will evaluate the data from the SNe Ia Pantheon+ sample (Brout et al., 2022) and the cosmic chronometers (CC) derived from the observational Hubble data compiled in Ref. (Moresco et al., 2022). We have not assumed that the Hubble and Pantheon datasets are correlated. Rather, we will present our results independently for the CC and Pantheon+ datasets. Utilising these datasets for statistical analysis enables us to obtain reliable results that are not influenced by assumptions of any particular underlying reference model (D’Agostino & Luongo, 2018; D’Agostino, 2019; Lohakare et al., 2023). In the following subsections, we will outline the key characteristics of these measurements and the corresponding likelihood functions.
3.1 Cosmic Chronometers (CC)
Figure 1: Contour plots show the $1-\sigma$ and $2-\sigma$ uncertainty regions for the variables $H_{0}$, $\Omega_{m0}$, $\alpha$, $\beta$, and $m$. These contours are derived from the CC sample (left panel) and the Pantheon+ data (right panel).
@@ 3.1 Cosmic Chronometers (CC)
$H(z)=\frac{\dot{a}}{a}=-\frac{1}{1+z}\frac{dz}{dt}\approx-\frac{1}{1+z}\frac{\Delta z}{\Delta t}\,,$ (14)
Here, $\dot{a}$ is the derivative of the scale factor $a$ with respect to time $t$, and $\Delta z$ and $\Delta t$ are the differences in redshift and time, respectively, between two objects. The value of $\Delta z$ can be determined by a spectroscopic survey, while the differential ages $\Delta t$ of passively evolving galaxies can be used to estimate the value of $H(z)$. Compiling such observations can be regarded as a CC sample. We use 32 objects spanning the redshift range $0.07\leq z\leq 1.965$ (Moresco et al., 2022). For these measurements, one can construct a $\chi^{2}_{\text{CC}}$ estimator as follows:
$\chi^{2}_{\text{CC}}=\sum_{i=1}^{32}\frac{[H_{\text{th}}(z_{i})-H_{\text{obs}}(z_{i})]^{2}}{\sigma^{2}_{H}(z_{i})}\,,$ (15)
@@
3.2 Supernovae type Ia (SNe Ia)
We will also take into account the Pantheon+ SNe Ia dataset, which includes 1701 measurements of the relative luminosity distance of SNe Ia spanning the redshift range of $2.26130.00122<z<$ 2.2613 (Brout et al., 2022). The Pantheon+ compilation consists of distance moduli derived from 1701 light curves of 1550 spectroscopically confirmed SNe Ia within the redshift range of $2.26130.00122<z<$ 2.2613, collected from 18 different surveys. It is worth noting that 77 of the 1701 light curves are associated with galaxies containing Cepheids. The Pantheon+ dataset is advantageous in that it can also be used to limit the value of $H_{0}$ besides the model parameters. To estimate the model parameter from the Pantheon+ samples, we minimize the $\chi^{2}$ function. To calculate the chi-square $(\chi^{2}_{\text{SNe}})$ value using the Pantheon compilation of 1701 supernovae data points, we use the following formula:
$\chi^{2}_{\text{SNe}}=\Delta\mu^{T}(C_{\text{Sys}+\text{Stat}}^{-1})\Delta\mu\,,$ (16)
The inverse covariance matrix, $C_{\text{Sys}+\text{Stat}}^{-1}$, associated with the Pantheon+ dataset, incorporates both systematic and statistical uncertainties. The term $\Delta\mu$, defined below, signifies the distance residual:
$\Delta\mu=\mu_{\text{th}}(z_{i},\theta)-\mu_{\text{obs}}(z_{i})\,.$ (17)
The distance modulus is specifically defined as the difference between an observed apparent magnitude ($m$) of the object and its absolute magnitude ($M$), which quantifies its intrinsic brightness. At a given redshift $z_{i}$, the distance modulus is expressed as follows:
$\mu_{\text{th}}(z_{i},\theta)=5\log_{10}\left(d_{L}(z,\theta)\right)+25=m-M\,,$ (18)
where $d_{L}$ denotes the luminosity distance in megaparsecs (Mpc), contingent upon the specific model, which is
$d_{L}(z,\theta)=\frac{c(1+z)}{H_{0}}\int_{0}^{z}\frac{d\zeta}{E(\zeta)}\,,$ (19)
where $E(z)=\frac{H(z)}{H_{0}}$, with $c$ representing the speed of light. Furthermore, the residual distance is indicated by
$\Delta\bar{\mu}=\begin{cases}\mu_{k}-\mu_{k}^{cd},&\text{if $k$ is in Cepheid hosts}\\ \mu_{k}-\mu_{\text{th}}(z_{k}),&\text{otherwise}\end{cases}$ (20)
where $\mu_{k}^{cd}$ represents the Cepheid host-galaxy distance as determined by SH0ES. This covariance matrix can be integrated with the SNe covariance matrix to form the covariance matrix for the Cepheid host galaxy. Incorporating both statistical and systematic uncertainties from the Pantheon+ dataset, the combined covariance matrix is expressed as $C^{\text{SNe}}_{\text{Sys}+\text{Stat}}+C^{cd}_{\text{Sys}+\text{Stat}}$. This formulation defines the $\chi^{2}$ function for the combined covariance matrix, which is utilized to constrain cosmological models in the analysis: